Find the area of the triangle having the given measurement. Round to the nearest square root
C= 42 degrees a= 4 feet b = 6 feet
A= 220 degrees b= 4 feet c= 5 feet

Answer by lwsshak3(11628)

Find the area of the triangle having the given measurement. Round to the nearest square root
C= 42 degrees a= 4 feet b = 6 feet
A= 220 degrees b= 4 feet c= 5 feet
***
Area of a triangle=(1/2)a*b*sinx,a and b are the two sides with the included angle x.
..
C=(1/2)*4*6*sin42˚=12*sin42˚=8.03 sq ft
A= (no solution) A<180˚in a triangle